At a gig 0.4 of the audience were female. If there were 100 more male audience members than femal, how many were there in the audience?
Answer by stanbon(75887)

At a gig 0.4 of the audience were female. If there were 100 more male audience members than female, how many were there in the audience?
-----
Equations:
f = 0.4(m + f)
m = f + 100
-----------------------
Rearrange:
0.6f = 0.4m
--------
Substitute for "f" and solve for "m"
0.6(m-100) = 0.4m
0.6m - 60 = 0.4m
0.2m = 60
m = 600/2 = 300 (# of males in the audience)
-------------
Cheers,
Stan H.